Segment AB has endpoints A(5,10) and B(-3,7). If the segment is extended 5 times its original length, what are the coordinates of the new end points?
Answer by Fombitz(32388)

I'm assuming that A stays fixed and that the new B moves since you didn't specify.
To get from A to B you move -3-5=-8 units in x and 7-10=-3 units in y.
If you multiply that by 5 you would get -40 units in x and -15 units in y.
Starting at A,
(5,10)+(-40,-15)=(-35,-5)
So the new B would be (-35,-5).